The measure of two complementary angles 16z - 9 and 4z + 3 > Find the measures of the angles.
please show the steps of how you solved this

Answer by jim_thompson5910(35256)

The two angles are complementary, so they must add to 90 degrees (by definition)


(Angle 1) + (Angle 2) = 90

(16z - 9) + ( 4z + 3 ) = 90

16z - 9 + 4z + 3 = 90

20z - 6 = 90

20z = 90+6

20z = 96

z = 96/20

z = 4.8

Now plug this back into the two original angle expressions
---------------------

Angle 1 = 16z - 9

Angle 1 = 16*4.8 - 9

Angle 1 = 67.8 degrees

---------------------

Angle 2 = 4z + 3

Angle 2 = 4*4.8 + 3

Angle 2 = 22.2 degrees